# Training Large Language Models to Reason In Parallel with Global Forking Tokens

### v2

@@ 2.3 Inference with Learned Global Forking Tokens
We discuss the inference protocols with $\mathrm{N}$ global forking tokens.
Figure 3: (Left) $\bm{\mathfrak{S}}_{p}$ is the subset of bipartite matching configurations that are still computed optimal towards the end of training. These can be individually visualized as a bipartite (Middle) Learned matchings by SSFT-32B in Exp 3, obtained by aggregating all edges in $\bm{\mathfrak{S}}_{p}$. (Right) At test time, for $\mathrm{Pass@}1$, we prompt with ${\textnormal{g}}^{(i^{\star})}$ that has the most connected edges. For $\mathrm{Cons@}k$, we prompt $i$-th parallel generation with <think(i % $\mathrm{N}$)>.
@@ 3 Experiments
We address the following research questions through experiments: (RQ1): In terms of $\mathrm{Pass@}1$ and $\mathrm{Cons@}k$ accuracy, how does a model trained with SSFT perform on reasoning benchmarks? (RQ2):  Does finding the optimal bipartite matching matter in reasoning performance? (RQ3): Does training with diverse reasoning traces yield better accuracy and coverage under SSFT compared to standard SFT with temperature scaling? (RQ4): Does prompting with distinct $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ genuinely make a model generate diverse reasoning traces? (RQ5): Is the performance gain from SSFT conditioned on the traces generated by our procedure and on a high-quality small dataset?
@@ 3.1 Experiment Setup
Training Dataset. We use the 1,000 questions from s1k dataset muennighoff2025s1_simpletesttime. In addition to the R1 (guo2025deepseekR1) and Gemini Flash (geminithinking) traces provided by s1, we also use Claude Opus 4.0/4.1 (claude_41_opus) and GPT-OSS-120B (agarwal2025gptoss) with high and medium reasoning effort to obtain a pool of distilled targets for the 1,000 questions. For each question, we generate two traces per source to populate the pool. We then sample four traces from this pool. We call this s1k-4mixed-reasoning dataset.
@@
Training Details. We fine-tune Qwen2.5-32B-Instruct (yang2025qwen3) for six epochs with a context length of 32,768. We reserve $\mathrm{N}=6$ global forking tokens and use $\mathrm{M}=4$ targets per question. To find the optimal bipartite matching for each input prompt, we consider only the first 1,000 tokens when computing the matching cost in Equation 4 for computational efficiency. We call this model SSFT-32B. We also include SSFT but choose a random bipartite matching at each step to fine-tune SSFT-32B (random $\bm{\sigma}$). Exact details on the pool of diverse distillation targets and selection procedure, as well as training hyperparameters, are provided in Appendix A.3.
@@
Baselines. All of our baselines use Qwen2.5-32B-Instruct as their base model, and only train on the 1k questions. Our baselines include two groups: (Single-Target $\triangle$) models trained with one trace per question and (Multi-Target $\bigstar$) models trained with four traces per question. For (Single-Target), we include s1.1-32B (muennighoff2025s1_simpletesttime), which uses 1k DeepSeek-R1 traces. We also fine-tuned an SFT-OSS-distill-32B baseline that trains only on the 1k GPT-OSS traces with high reasoning effort, as these traces achieved the highest correctness on the 1k questions based on an evaluation by Claude 3.5 Sonnet comparing each attempt against the reference answer. For (Multi-Target), we use our s1k-4mixed-reasoning to fine-tune SFT-mixed-distill-32B using standard SFT with one <think> token, duplicating each question and treating the four traces as four individual data points. We also include Multiverse-32B (yang2025multiverse), which prompts Gemini 2.5-Pro (Google:Gemini2.5Pro2025) to transform 1k sequential CoTs into parallel CoTs as their training data.
@@
Evaluation Setup. Our evaluation tasks consist of AIME24/AIME25 (ye2025aimepreview), MATH-500 (hendrycks2021measuringmath500), and GPQA-Diamond (rein2024gpqa). We use LightEval (lighteval) as our evaluation framework with generation configurations: temperature=0.7 used in (guha2025openthoughts), top_p=0.95, max length=32768. For $\mathrm{Pass@}1$ accuracy without any parallel test-time compute, we select learned ${\textnormal{g}}^{(1)}$ for SSFT-32B and ${\textnormal{g}}^{(4)}$ for SSFT-32B (random $\bm{\sigma}$) based on Equation 6. For each $\mathrm{Pass@}1$ accuracy, we compute the average performance over 32 generations. For $\mathrm{Cons@}6$, which applies each of the six global forking tokens once in our method and uses six generations for the baselines, we compute the average over 11 sets of generations to reduce variance in the results. We refer to this as Average of Native Cons@6, which corresponds to the Pass@1 metric with parallel generations in concurrent work (wen2025parathinker). Appendix  A.4 presents an example of the parallel generations by SSFT-32B.
@@ 3.2 Evaluating SSFT on Reasoning Benchmarks
AIME 2024 AIME 2025 MATH-500 GPQA-D Average
Pass@1: Average performance of individual generations
Qwen2.5-32B-Instruct $\triangle$ 15.80 10.40 80.40 47.00 38.40
s1.1-32B $\triangle$ 54.79 44.27 92.16 62.12 63.34
Multiverse-32B $\bigstar$ 53.80 45.80 91.80 60.70 63.03
@@
SSFT-32B (random $\bm{\sigma}$) $\bigstar$ 61.77 55.10 89.95 62.28 67.28
SSFT-32B $\bigstar$ 64.06 58.13 90.02 60.39 68.15
Average of Native Cons@6: Average performance of majority voting with 6 parallel generations
s1.1-32B $\triangle$ 70.30 53.33 95.60 61.45 70.17
@@
[…]
SSFT-32B (random $\bm{\sigma}$) $\bigstar$ 73.03 67.58 95.67 61.87 74.54
SSFT-32B $\bigstar$ 75.45 73.94 96.47 63.05 77.23
s1.1-32B $\triangle$ 73.33 63.33 94.80 60.61 73.02
[…]
SSFT-32B (random $\bm{\sigma}$) $\bigstar$ 80.00 80.00 95.60 62.63 79.56
SSFT-32B $\bigstar$ 83.33 86.67 96.80 61.62 82.11
@@
Table 1: Performance of SSFT compared to baselines on four reasoning tasks, reported at $\mathrm{Pass@}1$, $\mathrm{Cons@}6$, and $\mathrm{Cons@}32$. SSFT selects <think1> for $\mathrm{Pass@}1$ and replaces 6 generations with 6 generations prompted by distinct <thinki> for $\mathrm{Cons@}k$. We observe consistent improvements over (i) SFT-OSS-distill-32B, which uses the 1k OSS-high traces; (ii) SFT-mixed-distill-32B, which uses the four mixed traces but treats them as individual data; and (iii) SSFT-32B (random $\bm{\sigma}$), which trains using Equation 5 but with a randomly chosen $\bm{\sigma}$.
For RQ1, we see in Table 1 that SSFT delivers the best $\mathrm{Pass@}1$ accuracy, 64.06 on AIME24 and 58.13 on AIME25, outperforming SFT-mixed-32B by $8.33\%$ and $6.57\%$, respectively. We also observe consistent improvements on all four tasks under parallel test-time compute at two scales, $\mathrm{Cons@}6$ and $\mathrm{Cons@}32$, over SFT-mixed-32B, which was trained on the same reasoning traces. Some notable results are $\mathrm{Cons@}6=73.94\%,\mathrm{Cons@}32=86.67\%$ on AIME25. To answer RQ2, we observe consistent improvements over SSFT-32B (random $\bm{\sigma}$), with especially strong gains at $\mathrm{Cons@}6$ on AIME25, where effectiveness with few parallel generations is critical. As shown later in Figures 5 and  6, optimal bipartite matching is essential for preventing collapsing reasoning modes.
@@
For RQ3, we compare our method against SFT-mixed-32B under various $k$ in $\mathrm{Pass@}k$ accuracy with 32 generations to assess generation coverage. Figure 4 shows that SSFT achieves higher coverage across nearly all values of $k$. SFT-mixed-32B requires more allowed attempts and higher temperature to match the coverage of SSFT at the cost of lowering its $\mathrm{Pass@}1$ and $\mathrm{Cons@}6$ accuracy.
@@ 3.3 Evaluating Parallel Reasoning Diversity and Learned Matchings
Figure 5: (SSFT, optimal matching). Distribution of thinking-token counts and average performance on AIME24 (left) and AIME25 (right) prompted by a distinct $\texttt{<think1>},\dots,\texttt{<think6>}$. Each accuracy is averaged over 11 generations, shown above each whisker. We observe clear specialization of reasoning modes. $\texttt{<think1>},\dots,\texttt{<think4>}$ trigger longer traces, better suited to many AIME problems, and these surpass all the average accuracy under SFT without bipartite matching in Figure 6. <think5> and <think6> favor concise reasoning, which can be suboptimal for some AIME tasks, yet still raise $\mathrm{Cons@}6$ beyond all baselines by adding diversity. The consistency of these length distributions across AIME24 and AIME25 indicates the difference is not due to randomness, and that these think tags truly initiate distinct yet consistent reasoning modes.
Figure 6: (SSFT, random matching). Distribution of thinking-token counts and average performance on AIME24 (left) and AIME25 (right) prompted by a distinct $\texttt{<think1>},\dots,\texttt{<think6>}$. Each accuracy is averaged over 11 generations, shown above each whisker. Without optimal bipartite matching during SFT, the think tags initiate reasoning traces with similar length distributions and achieve similar average performance. Moreover, the slight length differences are not consistent across AIME24 and AIME25, suggesting randomness rather than genuine specialization; the tags therefore collapse to similar modes, failing to preserve the diverse reasoning modes in the training data. Both the average performance and $\mathrm{Cons@}6$ fall behind those from any think tag that initiates long reasoning under SSFT in Figure 5.
Addressing RQ4, we show that our global forking tokens genuinely initiate distinct reasoning traces and offer a new mechanism for leveraging test-time compute. We call these tokens emergent because we do not manually identify distinct reasoning modes over the traces for different questions. For example, <think1> could lead to longer reasoning than <think5> for one question, but this may not be consistent for other similar questions. We do not add any auxiliary objective to skew the original language modeling objective.
@@
Emerging Diverse Reasoning Modes. Using the $\mathrm{Cons@}6$ results in Table 1, we form six sets of generations, each prompted by a distinct ${\textnormal{g}}^{(i)}$. For each set, we show the average accuracy and the distribution of thinking-token counts: Figure 5 for SSFT with optimal bipartite matching and Figure 6 for SSFT with random matching. Reasoning length: Length partially indicates the diversity in reasoning, and we see clear differences for SSFT with optimal matching, despite the absence of hand-crafted matching rule or information about these traces. The consistency of these distributions across AIME24 and AIME25 indicates the differences is not from randomness, whereas randomly assigning a <thinki>, as in concurrent work (wen2025parathinker), does not yield clear or consistent differences in reasoning length, as shown in Figure 6. Performance: After finetuning with random matching, prompting with a distinct <thinki> shows no meaningful impact ($\approx 61\%$ on AIME24 and $\approx 55\%$ on AIME25). With optimal matching, SSFT elicits distinct reasoning modes initiated by $\texttt{<think1>},\dots,\texttt{<think4>}$ that reach around $65\%$ on AIME24 and $\geq 59\%$ on AIME25, with different lengths. Although <think5> and <think6> are weaker due to shorter reasoning modes, the average of these and, especially, $\mathrm{Cons@}6$ performance improve consistently with them.
@@
Figure 7: Cumulative counts of $\bm{\sigma}_{k}\in\bm{\mathfrak{S}}_{\mathrm{P}}$ computed as optimal over training. Note that $\bm{\mathfrak{S}}_{\mathrm{P}}$ and $\bm{\mathfrak{S}}_{p}$ are defined in Sections 2.1 and 2.3, respectively. Front axis: matching configuration index $k$. Depth: training step $t$. Bar height is the cumulative counts. These are the evolution of matchings during training 3 Qwen-32B-Instruct models under 3 bipartite matching settings. In this case study, the $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$ are always (GPT-OSS-high, GPT-OSS-med, R1, Gemini) for each question. Note that Random matching method does not minimize Eqn 5 under optimal matching, but we track it. We observe $\bm{\mathfrak{S}}_{p}=\bm{\mathfrak{S}}_{\mathrm{P}}$ with random matching, meaning no correlations learned. But by optimizing Hungarian loss, we see the emergence of $\bm{\mathfrak{S}}_{p}\subset\bm{\mathfrak{S}}_{\mathrm{P}}$.
(d) Visualization of learned matchings after SSFT with random matching. The fully connected graph indicates no correlations were learned.
(e) Visualization of learned matchings after SSFT with optimal matching between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{4}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$
(f) Visualization of learned matchings after SSFT with optimal matching between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{6}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$
Figure 8: We visualize the learned matchings by connecting edges in the subset of configurations $\bm{\mathfrak{S}}_{p}$ that still accumulate mass late in training (Figure 7). All the three models in (a), (b), and (c) are fine-tuned using GPT-OSS-high, GPT-OSS-medium, R1, and Gemini to distill reasoning traces for every question, enabling clearer interpretation.
Visualization of Learned Matchings. We observe in Figure 3 that some ${\textnormal{g}}^{(i)}\in{\bm{g}}$ from SSFT-32B have unique configuration of matched edges with $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$. This is a positive indication that $\{{\textnormal{g}}^{(i)}\}_{i=1}^{6}$ are likely to initiate distinct reasoning modes. We hypothesize that $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ can still yield a unique edge-matching configuration even if a subset of $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$ are difficult to distinguish.
For interpretability, we fine-tune using the same four teacher models for each $\mathbf{x}$}: GPT-OSS high, medium, R1, and Gemini. We call this dataset s1k-4teachers-reasoning dataset, and ask whether SSFT associates a unique ${\textnormal{g}}^{(i)}$ to the Gemini and R1 traces which have easily identifiable reasoning patterns, and then have the rest matched to OSS-high/OSS-medium traces in different ways (i.e. only matched to OSS-high, only matched to OSS-med, and matched to both). We track source indices only for evaluation; the model still receives an unordered set of traces with no source information for each question. We study three bipartite matching settings with SSFT. Figure 7 shows the evolution of matchings learned under these three hyperparameters: (a) random matching with four reserved ${\textnormal{g}}^{(i)}$, (b) optimal matching with four reserved ${\textnormal{g}}^{(i)}$, and (c) optimal matching with six reserved ${\textnormal{g}}^{(i)}$. Initially, all the configurations $\bm{\sigma}\in\bm{\mathfrak{S}}_{\mathrm{P}}$ accumulate mass as there is no correlations between $\bm{g}$} sand $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$. Figure 7(a) shows SSFT under random matching does not shrink the size of configurations computed as optimal, meaning that no correlations are learned between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{4}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$. By contrast, Figure 7(b) and Figure 7(c) show the emergence of only a strict subset of matching configurations in $\bm{\mathfrak{S}}_{\mathrm{P}}$. This indicates some correlations between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ are indeed learned through SSFT. We first visualize the learned matchings of the model with four ${\textnormal{g}}^{(i)}$ in Figure 8. We observe that ${\textnormal{g}}^{(1)}$ and ${\textnormal{g}}^{(2)}$ are uniquely matched to the R1 and Gemini traces, showing that SSFT can indeed uniquely associate ${\textnormal{g}}^{(i)}$ to sufficiently diverse reasoning traces. Now to confirm our previous hypothesis, we see the unique learned matchings $({\textnormal{g}}^{(3)},(\text{OSS-high, OSS-med})),({\textnormal{g}}^{(4)},\text{OSS-med})$. Furthermore, by connecting all the edges in $\bm{\mathfrak{S}}_{p}$ from SSFT with 6 forking tokens (Figure 7(c)), we also see unique learned matchings $({\textnormal{g}}^{(3)},\text{OSS-high}),({\textnormal{g}}^{(4)},\text{OSS-med}),({\textnormal{g}}^{(5)},(\text{OSS-med},\text{OSS-high}))$ in Figure 8. This confirms that the global forking tokens can identify unique correlations even among highly similar traces.
3.4 Ablation Study: Removing High Quality Small Dataset

### v3

@@ 2.4 Inference with Learned Global Forking Tokens
Figure 2: Learned matchings by SSFT-32B in Exp 3, obtained by connecting edges in $\bm{\mathfrak{S}}_{p}$.
@@ 3 Experiments
We address the following research questions through experiments: (RQ1): In terms of $\mathrm{Pass@}1$ and $\mathrm{Cons@}k$ accuracy, how does a model trained with SSFT perform on reasoning benchmarks? (RQ2):  Does finding the optimal bipartite matching matter in reasoning performance? (RQ3): Does training with diverse reasoning traces yield better accuracy and coverage under SSFT compared to standard SFT with temperature scaling? (RQ4): Does prompting with distinct $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ genuinely make a model generate diverse reasoning traces? (RQ5): Is the performance gain from SSFT robust to different fine-tuning datasets and model scales?
@@ 3.1 Experiment Setup
Training Dataset. We use the 1,000 questions from s1k dataset Muennighoff et al. (2025). In addition to the R1 (Guo et al., 2025) and Gemini Flash (Google, 2024) traces provided by s1, we also use Claude Opus 4.0/4.1 (Anthropic, 2025) and GPT-OSS-120B (Agarwal et al., 2025) with high and medium reasoning effort to obtain a pool of distilled targets for the 1,000 questions. For each question, we generate two traces per source to populate the pool. We then sample four traces from this pool. We call this s1k-4mixed-reasoning dataset. For additional GFPO, we use 1280 questions from DAPO-Math-17k (Yu et al., 2025).
@@
Training Details. We fine-tune Qwen2.5-32B-Instruct (Yang et al., 2025a) for six epochs with a context length of 32,768. We reserve $\mathrm{N}=6$ global forking tokens and use $\mathrm{M}=4$ targets per question. To find the optimal bipartite matching for each input prompt, we consider only the first 1,000 tokens when computing the matching cost in Equation 2. This number is derived from $\lfloor 32768/(\mathrm{M}\mathrm{N})\rfloor$ for computational efficiency, so all matching costs fit in one forward pass without extra VRAM. We call this model SSFT-32B. We also include SSFT but choose a random bipartite matching at each step to fine-tune SSFT-32B (random $\bm{\sigma}$). Models with an additional RL step use the postfix GRPO/GFPO. Exact details on the pool of diverse distillation targets and selection procedure, as well as training hyperparameters, are provided in Appendix A.9 and Table  9.
@@
Baselines. Our baselines include two groups: (Single-Target $\triangle$) models trained with one trace per question and (Multi-Target $\bigstar$) models trained with four traces per question. For (Single-Target), we include s1.1-32B (Muennighoff et al., 2025), which uses 1k DeepSeek-R1 traces. We also fine-tuned an SFT-OSS-distill-32B baseline that trains only on the 1k GPT-OSS traces with high reasoning effort, as these traces achieved the highest correctness on the 1k questions based on an evaluation by Claude 3.5 Sonnet comparing each attempt against the reference answer. For (Multi-Target), we use s1k-4mixed-reasoning to fine-tune SFT-mixed-distill-32B-tags using standard SFT, where each question is duplicated for its distilled traces and <think i>  is sampled without replacement for each duplicate, treating the four traces as four individual data points. We also include Multiverse-32B (Yang et al., 2025b), which prompts Gemini 2.5-Pro (Google, 2025a) to transform 1k sequential CoTs into parallel CoTs as their training data.
@@
Evaluation Setup. Our evaluation tasks consist of AIME24/AIME25 (Ye et al., 2025), MATH-500 (Hendrycks et al., 2021), GPQA-Diamond (Rein et al., 2024), and LiveCodeBench (LCB), where LCB is considered out-of-distribution relative to the fine-tuning dataset. We use LightEval (Habib et al., 2023) as our evaluation framework with generation configurations: temperature=0.7 used in (Guha et al., 2025), top_p=0.95, max length=32768. For $\mathrm{Pass@}1$ accuracy without any parallel test-time compute, we select learned ${\textnormal{g}}^{(1)}$ for SSFT-32B and ${\textnormal{g}}^{(4)}$ for SSFT-32B (random $\bm{\sigma}$) based on Equation 4. We let the model sample ${\textnormal{g}}^{(i)}$ for SSFT-32B-GFPO. For each $\mathrm{Pass@}1$ accuracy, we compute the average performance over 32 generations. For $\mathrm{Cons@}6$, which applies each of the six global forking tokens once in our method and uses six generations for the baselines, we compute the average over 11 sets of generations to reduce variance in the results. We refer to this as Pass@1 of Native Cons@6 using a similar terminology as the concurrent work (Wen et al., 2025). Appendix  A.10 presents an example of the parallel generations.
@@ 3.2 Evaluating SSFT on Reasoning Benchmarks
Table 1: Performance of SSFT compared to baselines on reasoning tasks, reported at $\mathrm{Pass@}1$, $\mathrm{Cons@}6$, and $\mathrm{Cons@}32$. SSFT selects <think 1> for $\mathrm{Pass@}1$ and replaces 6 generations with 6 generations prompted by distinct <think i> for $\mathrm{Cons@}k$. SSFT-GFPO samples <think i> for $\mathrm{Pass@}1$. We observe consistent improvements over (i) SFT-OSS-distill-32B, which uses the 1k OSS-high traces; (ii) SFT-mixed-distill-32B-tags, which uses four mixed traces but treats them as individual data; and (iii) SSFT-32B (random $\bm{\sigma}$), which optimizes Equation 3 with a random $\bm{\sigma}$.
AIME 2024 AIME 2025 MATH-500 GPQA-D $\text{LCB(v5)}^{*}$
Pass@1: Average performance of individual generations (* denotes out-of-distribution tasks.)
Qwen2.5-32B-Instruct $\triangle$ 15.80 10.40 80.40 47.00 23.35
s1.1-32B $\triangle$ 54.79 44.27 92.16 62.12 -
Multiverse-32B $\bigstar$ 53.80 45.80 91.80 60.70 -
@@
SFT-mixed-distill-32B-tags $\bigstar$ 58.23 51.96 88.49 59.96 32.34
SSFT-32B (random $\bm{\sigma}$) $\bigstar$ 61.77 55.10 89.95 62.28 35.33
SSFT-32B $\bigstar$ 64.06 58.13 90.02 60.39 38.92
SSFT-32B-GFPO $\bigstar$ 64.22 58.80 89.90 62.48 42.10
Pass@1 of Native Cons@6: Average performance of majority voting with 6 parallel generations
s1.1-32B $\triangle$ 70.30 53.33 95.60 61.45 -
@@
SFT-mixed-distill-32B-tags $\bigstar$ 73.94 70.00 95.88 58.75 -
SSFT-32B (random $\bm{\sigma}$) $\bigstar$ 73.03 67.58 95.67 61.87 -
SSFT-32B $\bigstar$ 75.45 73.94 96.47 63.05 -
SSFT-32B-GFPO $\bigstar$ 76.67 78.48 95.73 60.94 -
s1.1-32B $\triangle$ 73.33 63.33 94.80 60.61 -
@@
SFT-mixed-distill-32B-tags $\bigstar$ 76.67 76.67 96.20 58.59
SSFT-32B (random $\bm{\sigma}$) $\bigstar$ 80.00 80.00 95.60 62.63 -
SSFT-32B $\bigstar$ 83.33 86.67 96.80 61.62 -
SSFT-32B-GFPO $\bigstar$ 83.33 83.33 96.80 62.12 -
@@
For RQ1, we see in Table 1 that SSFT delivers the best $\mathrm{Pass@}1$ accuracy, 64.06 on AIME24 and 58.13 on AIME25, outperforming SFT-mixed-32B by $8.33\%$ and $6.57\%$, respectively. We also observe consistent improvements on all four tasks under parallel test-time compute at two scales, $\mathrm{Cons@}6$ and $\mathrm{Cons@}32$, over SFT-mixed-32B, which was trained on the same reasoning traces. Some notable results are $\mathrm{Cons@}6=73.94\%,\mathrm{Cons@}32=86.67\%$ on AIME25. To answer RQ2, we observe consistent improvements over SSFT-32B (random $\bm{\sigma}$), with especially strong gains at $\mathrm{Cons@}6$ on AIME25, where effectiveness with few parallel generations is critical. As shown later in Figures 4 and  5, optimal bipartite matching is essential to prevent collapsing reasoning modes. Results with SSFT-GFPO confirm that GFPO successfully sharpens the output distribution of ${\textnormal{g}}^{(i)}$ conditioned on $\mathbf{x}$}. We even observe positive gains in $\mathrm{Pass@}1$. Appendix A.3 reports the results for baselines with RL. For out-of-distribution generalization to LCB, SSFT also yields the largest gains. Coverage. For RQ3, we compare our method against SFT-mixed-32B under various $k$ in $\mathrm{Pass@}k$ accuracy with 32 generations to assess generation coverage. Figure 3 shows that SSFT achieves higher coverage across nearly all values of $k$. SFT-mixed-32B requires more allowed attempts and higher temperature to match the coverage of SSFT at the cost of lowering its $\mathrm{Pass@}1$ and $\mathrm{Cons@}6$.
@@ 3.3 Evaluating Parallel Reasoning Diversity and Learned Matchings
Figure 4: (SSFT, optimal matching). Distribution of thinking-token counts and average performance on AIME24 (left) and AIME25 (right) prompted by a distinct $\texttt{<think\,1>},\dots,\texttt{<think\,6>}$.
Figure 5: (SSFT, random matching). Distribution of thinking-token counts and average performance on AIME24 (left) and AIME25 (right) prompted by a distinct $\texttt{<think\,1>},\dots,\texttt{<think\,6>}$.
Addressing RQ4, we show that our global forking tokens genuinely initiate distinct reasoning traces and offer a new mechanism for leveraging test-time compute.
@@
Emerging Diverse Reasoning Modes. Using the $\mathrm{Cons@}6$ results in Table 1, we form six sets of generations, each prompted by a distinct ${\textnormal{g}}^{(i)}$. For each set, we show the average accuracy and the distribution of thinking-token counts: Figure 4 for SSFT with optimal bipartite matching and Figure 5 for SSFT with random matching. Reasoning length: Length partially indicates the diversity in reasoning, and we see clear differences for SSFT with optimal matching, despite the absence of hand-crafted matching rule or information about these traces. The consistency of these distributions across AIME24 and AIME25 indicates the differences is not from randomness, whereas randomly assigning a <think i>, as in concurrent work (Wen et al., 2025), does not yield clear or consistent differences in reasoning length, as shown in Figure 5. Performance: After finetuning with random matching, prompting with a distinct <think i> shows no meaningful impact ($\approx 61\%$ on AIME24 and $\approx 55\%$ on AIME25). With optimal matching, SSFT elicits distinct reasoning modes initiated by $\texttt{<think\,1>},\dots,\texttt{<think\,4>}$ that reach around $65\%$ on AIME24 and $\geq 59\%$ on AIME25, with different lengths. Although <think 5> and <think 6> are weaker due to shorter reasoning modes, the average of these and, especially, $\mathrm{Cons@}6$ performance improve consistently with them.
Visualization of Learned Matchings. We observe in Figure 2 that some ${\textnormal{g}}^{(i)}\in{\bm{g}}$ from SSFT-32B have unique configuration of matched edges with $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$. This is a positive indication that $\{{\textnormal{g}}^{(i)}\}_{i=1}^{6}$ are likely to initiate distinct reasoning modes. We hypothesize that $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ can still yield a unique edge-matching configuration even if a subset of $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$ are difficult to distinguish. For interpretability, we fine-tune using the same four teacher models for each $\mathbf{x}$}: GPT-OSS high, medium, R1, and Gemini. We call this dataset s1k-4teachers-reasoning dataset, and ask whether SSFT associate a unique ${\textnormal{g}}^{(i)}$ to the Gemini and R1 traces which have easily identifiable reasoning patterns, and then have the rest matched to OSS-high/OSS-medium traces in different ways (i.e. only matched to OSS-high, only matched to OSS-med, and matched to both). We track source indices only for evaluation; the model still receives an unordered set of traces with no source information for each question. We study 3 bipartite matching settings for SSFT. Figure 6 shows the evolution of matchings learned under these 3 hyperparameters SSFT (a) random matching with four reserved ${\textnormal{g}}^{(i)}$, (b) optimal matching with four reserved ${\textnormal{g}}^{(i)}$, and (c) optimal matching with six reserved ${\textnormal{g}}^{(i)}$. Initially, all the configurations $\bm{\sigma}\in\bm{\mathfrak{S}}_{\mathrm{P}}$ accumulate mass as there is no correlations between $\bm{g}$} sand $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$. Figure 6(a) shows SSFT under random matching does not shrink the size of configs computed as optimal, meaning that no correlations are learned between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{4}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$. By contrast, Figure 6(b) and Figure 6(c) show the emergence of only a strict subset of matching configurations in $\bm{\mathfrak{S}}_{\mathrm{P}}$. This indicates some correlations between $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ are indeed learned through SSFT. We first visualize the learned matchings of the model with four ${\textnormal{g}}^{(i)}$ in Figure 8. We observe that ${\textnormal{g}}^{(1)}$ and ${\textnormal{g}}^{(2)}$ are uniquely matched to the R1 and Gemini traces, showing that SSFT can indeed uniquely associate ${\textnormal{g}}^{(i)}$ to sufficiently diverse reasoning traces. Now to confirm our previous hypothesis, we see the unique learned matchings $({\textnormal{g}}^{(3)},(\text{OSS-high, OSS-med})),({\textnormal{g}}^{(4)},\text{OSS-med})$. Furthermore, by connecting all the edges in $\bm{\mathfrak{S}}_{p}$ from SSFT with 6 forking tokens (Figure 6(c)), we also see unique learned matchings $({\textnormal{g}}^{(3)},\text{OSS-high}),({\textnormal{g}}^{(4)},\text{OSS-med}),({\textnormal{g}}^{(5)},(\text{OSS-med},\text{OSS-high}))$ in Figure 8. This confirms that the global forking tokens can identify unique correlations even among highly similar traces.
@@
Figure 6: Cumulative counts of $\bm{\sigma}_{k}\in\bm{\mathfrak{S}}_{\mathrm{P}}$ computed as optimal over training. Note that $\bm{\mathfrak{S}}_{\mathrm{P}}$ and $\bm{\mathfrak{S}}_{p}$ are defined in Sections 2.1 and 2.4, respectively. Front axis: matching configuration index $k$. Depth: training step $t$. Bar height is the cumulative counts. These are the evolution of matchings during training 3 Qwen-32B-Instruct models under 3 bipartite matching settings. In this case study, the $\{{\mathbf{r}}^{(j)}\}_{j=1}^{4}$ are always (GPT-OSS-high, GPT-OSS-med, R1, Gemini) for each question. Note that Random matching method does not minimize Eqn 3 under optimal matching, but we track it. We observe $\bm{\mathfrak{S}}_{p}=\bm{\mathfrak{S}}_{\mathrm{P}}$ with random matching, meaning no correlations learned. But by optimizing Hungarian loss, we see the emergence of $\bm{\mathfrak{S}}_{p}\subset\bm{\mathfrak{S}}_{\mathrm{P}}$.
4 Robustness and Ablation Studies
Table 2: Comparison of SSFT and baseline fine-tuned models trained on code-generation data, evaluated on one coding task and three math reasoning tasks that are out-of-distribution relative to the fine-tuning set. At test time, <think i>  is sampled by the models.
LCB(v5) $\text{AIME24}^{*}$ $\text{AIME25}^{*}$ $\text{MATH500}^{*}$
Pass@1 w/ 16 samples for LCB. (* denotes OOD w.r.t post-training)
Qwen2.5-32B-Instruct 23.35 15.80 10.40 80.40
SFT-mixed-distill-32B-code 47.13 34.69 24.17 89.39
SSFT-32B-code (random) 45.36 39.06 32.92 89.46
SSFT-32B-code 52.07 43.23 32.82 89.96
Robustness to Fine-tuning with Code Generation traces.   We study whether SSFT can improve $\mathrm{Pass@}1$ on more open-ended tasks by instilling specialized, complex reasoning modes without mode collapse, unlike standard SFT where such modes can be corrupted by merging together. We select 1K coding questions from Open-Thoughts (Guha et al., 2025), and use R1 to generate four reasoning traces per question to construct code1k-4mixed-reasoning. We apply SSFT on this dataset, and then GFPO on 1k questions from Intellect-2-RL (Team et al., 2025). From the $\mathrm{Pass@}1$ comparison in Table 2, SSFT better elicits complex reasoning modes of higher accuracy than baselines that do not use a set loss. Full results are in Appendix 5.